### v2

@@ Appendix A Opial’s lemma for quasi-Féjer monotonicity
Here we prove a generalisation of Opial’s lemma [23] for quasi-Féjer monotonicity, i.e, Féjer monotonicity with an additive error term. We prove it in normed spaces for Bregman divergences (47), as they add no extra difficulties. In an even more general variable-metric framework, a similar result is also proved in [21, Proposition 2.7]. Our simplified proof follows the outline of that in [7], and is nearly identical to the one in [34], where the errors took a more specific form.
@@
Proof A.3.
Let $\bar{x}$ and $\hat{x}$ be weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$. Since Bregman divergences $B_{M}\geq 0$ for convex $M$, the conditions Items (iii) and (ii) establish the assumptions of Lemma A.1 for $a_{k}=B_{M}(x^{k};\bar{x})$, $b_{k}=0$, $c_{k}=e_{k}(\bar{x})$, and $d_{k}=0$. It follows that $\{B_{M}(x^{k};\bar{x})\}_{k\in\mathbb{N}}$ is convergent. Likewise we establish that $\{B_{M}(x^{k};\hat{x})\}_{k\in\mathbb{N}}$ is convergent. Therefore, by the obvious three-point identity for Bregman divergences (see, e.g., [32]),
@@
If $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded, and $X$ is the dual space of some separable normed space $X_{*}$, it contains a weakly-* convergent subsequence by the Banach–Alaoglu theorem, so a limit point exists as claimed. If $X$ is reflexive, the Eberlein–S̆mulyan theorem establishes the same result. Hence, if (57) implies $\bar{x}=\hat{x}$, then every convergent subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ has the same weak limit. It lies in $\hat{X}$ by Item (i). The final claim now follows from a standard subsequence–subsequence argument: Assume to the contrary that there exists a subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ not convergent to $\hat{x}$. Then the above argument provides a further subsequence converging to $\hat{x}$. This contradicts the fact that any subsequence of a convergent sequence converges to the same limit.

### v3

@@ Appendix B Opial’s lemma for quasi-Féjer monotonicity
Here we prove a generalisation of Opial’s lemma [25] for quasi-Féjer monotonicity, i.e, Féjer monotonicity with an additive error term. We prove it in normed spaces for Bregman divergences (34), as they add no extra difficulties. In an even more general variable-metric framework, a similar result is also proved in [23, Proposition 2.7]. Our simplified proof follows the outline of that in [7], and is nearly identical to the one in [35], where the errors took a more specific form.
@@
Proof B.3.
Let $\bar{x}$ and $\hat{x}$ be weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$. Since Bregman divergences $B_{M}\geq 0$ for convex $M$, the conditions Items˜(iii) and (ii) establish the assumptions of Lemma˜B.1 for $a_{k}=B_{M}(x^{k};\bar{x})$, $b_{k}=0$, $c_{k}=e_{k}(\bar{x})$, and $d_{k}=0$. It follows that $\{B_{M}(x^{k};\bar{x})\}_{k\in\mathbb{N}}$ is convergent. Likewise we establish that $\{B_{M}(x^{k};\hat{x})\}_{k\in\mathbb{N}}$ is convergent. Therefore, by the obvious three-point identity for Bregman divergences (see, e.g., [33]),
@@
If $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded, and $X$ is the dual space of some separable normed space $X_{*}$, it contains a weakly-* convergent subsequence by the Banach–Alaoglu theorem, so a limit point exists as claimed. If $X$ is reflexive, the Eberlein–S̆mulyan theorem establishes the same result. Hence, if (75) implies $\bar{x}=\hat{x}$, then every convergent subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ has the same weak limit. It lies in $\hat{X}$ by Item˜(i). The final claim now follows from a standard subsequence–subsequence argument: Assume to the contrary that there exists a subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ not convergent to $\hat{x}$. Then the above argument provides a further subsequence converging to $\hat{x}$. This contradicts the fact that any subsequence of a convergent sequence converges to the same limit.